OpenStudy (anonymous):

Solve the equation:2x-8=-8(3x-1/2)

OpenStudy (anonymous):

2x-8=-8(3x-1/2) 2x-8 = -24x+4 // add 24x to both sides. add 8 to both sides. solve for x

OpenStudy (anonymous):

i mostly get stuck when dealing with fractions

OpenStudy (anonymous):

ok, so we have 2x-8= -24x+4 2x+24x=4+8 26x=12 x = 12/26 = 6/13

OpenStudy (anonymous):

so when you distribute a fraction you just multiply straight across? like 8/1x1/2

OpenStudy (anonymous):

yeah, so it becomes 8/2 or 4